find the average velocity of a particle released from rest from a hieght of 125 m over time interval till it sirkes the ground
 
 

Grade:12

1 Answers

Susmita
5 years ago
At height h=125m,velocity u=0.
h=ut+gt2/2=gt2/2=5t2 taking g=10m/s2
or,t=root(h/5)=root(125/5)=5s
So average speed=(distance covered)/time elasped=125/5=25 m/s
Hope this helps.